You roll a dice. If the number is even you lose that amount of money, if it is odd you win that amount. Let W be your winnings after one turn, what is EW?
Answer #1

Let X shows the amount the money you lose when you get even number and Y shows the amount of monet you lose.

When a fair die is rolled possible outcoes are 1, 2, 3, 4, 5, 6. Out of 6 outcome 3 are even and 3 are odd. So

P(even) = 3/6 =0.5

P(odd) = 3/6 = 0.5

Here X can take values X and Y. So expected value of W is

E(W) = P(even) * X + P(odd) * Y = 0.5X +0.5Y = 0.5 (X+Y)

Hence, E(W) = 0.5 (X+Y)